OpenStudy (anonymous):

line AB = BC with angle 1= 6x-13 and angle 3= 3x +15 find the measure of angle 2

OpenStudy (anonymous):

94

OpenStudy (anonymous):

angle 1 = angle3 6x-13=3x+15 find x x=28/3 find angle1 and angle 2 sum of three angles is 180 find the third angle

OpenStudy (anonymous):

options are 19 25 35 and 43

OpenStudy (anonymous):

a what angle ?

OpenStudy (anonymous):

vertical angles one makes a triangle